### v1

# Integrable open elliptic Toda chain with boundaries

Abstract
In this letter we discuss the classical integrable elliptic Toda chain proposed by I. Krichever. Our goal is to construct an open elliptic Toda chain with boundary terms. This is achieved using the factorized form of the Lax matrix and gauge equivalence with the XYZ chain.
1 Introduction
The elliptic closed Toda chain was introduced by I. Krichever [6]. It is a classical integrable system defined by the following Hamiltonian:
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{eToda}}}=-\frac{1}{2}\sum\limits_{a=1}^{n}\Big(\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\Big)\,,}\end{array}$ (1)
where ${\bf{p}}_{a}$, ${\bf{q}}_{a}$, a=1,,na=1,...,n is a set of canonically conjugated variables (that is $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$) and $\wp(x)$ is the Weierstrass elliptic function on elliptic curve with moduli $\tau$, ${\rm Im}(\tau)>0$. We assume here that numeration of particles is modulo $n$: ${\bf{q}}_{n+1}={\bf{q}}_{1}$, ${\bf{q}}_{0}={\bf{q}}_{n}$. This model is a particular case of a more general elliptic Ruijsenaars-Toda chain [2], see also [1, 14, 15].
It was recently shown in [10] that the elliptic Toda chain is a particular case of the Ruijsenaars chain introduced previously in [16]. The classical $r$-matrix structure for the Ruijsenaars chain was derived in [9] and then adopted to the elliptic Toda chain in [10]. Another important observation of [10] and [16] is that the elliptic Toda chain is gauge equivalent to the discrete Landau-Lifshitz model (or the classical XYZ chain) with some special restrictions. In the next two sections we briefly recall the main results on this topic.
The purpose of the paper is to derive an open elliptic Toda chain with boundary terms using the gauge equivalence mentioned above. Our construction is based on the standard description of the classical transfer-matrix for XYZ chain with boundary [13]. Using the factorized form for the Lax matrices we represent the transfer-matrix as the one describing the open Toda chain with the gauge transformed $K$-matrices. This allows to compute the Hamiltonian, and the integrability follows from integrability of the XYZ chain with boundaries.
2 Lax matrices in factorized form
Introduce the following matrix:
$\begin{array}[]{c}\displaystyle{g(z,{\bf{q}}_{a})=\frac{1}{\theta_{1}(2{\bf{q}}_{a}|\tau)}\left(\begin{array}[]{cc}{\theta_{3}(z-2{\bf{q}}_{a}|2\tau)}&{-\theta_{3}(z+2{\bf{q}}_{a}|2\tau)}\\ \\ {-\theta_{2}(z-2{\bf{q}}_{a}|2\tau)}&{\theta_{2}(z+2{\bf{q}}_{a}|2\tau)}\end{array}\right)\,.}\end{array}$ (2)
This is an intertwining matrix entering the IRF-Vertex correspondence [3] in quantum statistical models. We use the standard notations for Jacobi theta functions. From here on it is assumed that all elliptic functions depend on $\tau$ unless otherwise stated explicitly (like in (2)). Let us introduce a notation for a set of $2\times 2$ matrices:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=-g^{-1}(z,{\bf{q}}_{k})g(z,{\bf{q}}_{m})e^{P^{m}/2c}\,,\qquad P^{m}={\rm diag}({\bf{p}}_{m},-{\bf{p}}_{m})}\end{array}$ (3)
Direct calculation yields expression for $L^{[k,m]}(z)$ in terms of the elliptic Kronecker function $\phi(z,x)$:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\\ {\phi(z,-{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,-{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\end{array}\right)\,,\quad\phi(z,u)=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+u)}{\theta_{1}(z)\theta_{1}(u)}}\end{array}$ (4)
where
$\begin{array}[]{c}\displaystyle{b^{[k,m]}_{1}=\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{{{\bf{p}}_{m}}/{2c}}\ \ b^{[k,m]}_{2}=-\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{-{{\bf{p}}_{m}}/{2c}}\,.}\end{array}$ (5)
The Lax matrices for the elliptic Toda chain can be written in several different ways. First possibility is to define the monodromy matrix $T^{\hbox{\tiny{eToda}}}(z)$ as
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=L^{[n,1]}(z)L^{[1,2]}(z)L^{[2,3]}(z)\dots L^{[n-1,n]}(z)\,,}\end{array}$ (6)
where we use the Lax matrices (4). Another possibility is to use the modified Lax representation [16, 10]:
$\begin{array}[]{c}\displaystyle{{\bf T}^{\hbox{\tiny{eToda}}}(z)={\bf L}^{1}(z){\bf L}^{2}(z)\dots\bf L^{n}(z)\,,\qquad{\bf L}^{a}(z)=\frac{1}{b^{[a-1,a]}_{1}+b^{[a-1,a]}_{2}}\,L^{[a-1,a]}(z)\,.}\end{array}$ (7)
Then ${\bf T}^{\hbox{\tiny{eToda}}}(z)={T}^{\hbox{\tiny{eToda}}}(z)e^{-H^{\hbox{\tiny{eToda}}}/2}$. Explicitly,
$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{2}\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\\ {\phi(z,-{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,-{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\end{array}\right)\,.}\end{array}$ (8)
Using the identity $\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)$ we have
$\begin{array}[]{c}\displaystyle{\det{\bf L}^{a}(z)=\frac{1}{4}\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (9)
Then it is easy to see that $\log\det{\bf T}^{\hbox{\tiny{eToda}}}(z)$ provides the Hamiltonian (1).
3 Gauge equivalence with XYZ chain
Plugging the factorized representation (3) into (6) gives
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=(-1)^{n}g^{-1}(z,{\bf{q}}_{n})g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})g(z,{\bf{q}}_{2})e^{P^{2}/2c}...g^{-1}(z,{\bf{q}}_{n-1})g(z,{\bf{q}}_{n})e^{P^{n}/2c}}\end{array}$ (10)
Therefore, the gauge transformed monodromy matrix ${\mathcal{T}}(z)=g(z,{\bf{q}}_{n})T^{\hbox{\tiny{eToda}}}(z)g^{-1}(z,{\bf{q}}_{n})$ takes the form
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=(-1)^{n}{\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)\,,\qquad{\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})}\end{array}$ (11)
It was explained in [10] that ${\mathcal{L}}^{a}(z)$ are the Lax matrices of the XYZ chain, and ${\mathcal{T}}(z)$ is its standard monodromy matrix [12, 4]. Namely,
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\,,\qquad\varphi_{k}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z)}{\theta_{1}(z)\theta_{5-k}(0)}\,,\quad k=1,2,3\,,}\end{array}$ (12)
where $\sigma_{a}$ are the Pauli matrices ($\sigma_{0}$ is $2\times 2$ identity matrix). The variables ${\bf S}^{a}_{k}$ are expressed through ${\bf{p}}_{a}$ and ${\bf{q}}_{a}$ as
$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}=\frac{1}{2}\Big(e^{{\bf{p}}_{a}/2c}+e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad{\bf S}_{k}^{a}=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)}\end{array}$ (13)
for all a=1,,na=1,...,n and $k=1,2,3$. The Poisson brackets between ${\bf S}_{k}^{a}$ computed through $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$ reproduce the classical Sklyanin algebra [11]
$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i},{\bf S}^{a}_{j}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}{\bf S}^{a}_{k}\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0},{\bf S}^{a}_{i}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}{\bf S}^{a}_{k}\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,,}\end{array}$ (14)
where the half-periods $\omega_{j}$
$\begin{array}[]{c}\displaystyle{\omega_{1}=\frac{\tau}{2}\,,\quad\omega_{2}=\frac{1+\tau}{2}\,,\quad\omega_{3}=\frac{1}{2}}\end{array}$ (15)
are numerated according to numeration of the Pauli matrices. The Sklyanin algebra (14) is generated by the classical quadratic exchange relation
$\begin{array}[]{c}\displaystyle{\{{\mathcal{L}}_{1}^{a}(z),{\mathcal{L}}_{2}^{b}(w)\}=\frac{\delta^{ab}}{c}\,[{\mathcal{L}}_{1}^{a}(z){\mathcal{L}}_{2}^{a}(w),r_{12}(z-w)]\,,\qquad r_{12}(z-w)=\frac{1}{2}\sum\limits_{k=1}^{3}\varphi_{k}(z-w)\sigma_{k}\otimes\sigma_{k}\,.}\end{array}$ (16)
The representation (13) for the variables ${\bf S}^{a}_{k}$ corresponds to special values of the Casimir functions ${\bf C}_{1}^{a}=({\bf S}^{a}_{1})^{2}+({\bf S}^{a}_{2})^{2}+({\bf S}^{a}_{3})^{2}$ and ${\bf C}_{2}^{a}=({\bf S}^{a}_{0})^{2}+\sum\limits_{k=1}^{3}({\bf S}^{a}_{k})^{2}\wp(\omega_{k})$:
$\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}=0\,,\qquad{\bf C}_{2}^{a}=1\,.}\end{array}$ (17)
4 XYZ chain with boundaries
Following [13] consider the monodromy matrix:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z){\mathcal{T}}(z)K^{-}(z){\mathcal{T}}^{-1}(-z)\,,}\end{array}$ (18)
where ${\mathcal{T}}(z)$ (11) satisfies the quadratic classical exchange relation
$\begin{array}[]{c}\displaystyle{c\{{\mathcal{T}}_{1}(z),{\mathcal{T}}_{2}(w)\}=[{\mathcal{T}}_{1}(z){\mathcal{T}}_{2}(w),r_{12}(z-w)]}\end{array}$ (19)
due to (16), and $K^{\pm}(z)$ are $2\times 2$ matrices solving the classical quadratic reflection equation
${[K^{\pm}_{1}(z)K^{\pm}_{2}(w),r_{12}(z-w)]+\,K^{\pm}_{2}(w)r_{12}(z+w)K^{\pm}_{1}(z)-\,K^{\pm}_{1}(z)r_{12}(z+w)K^{\pm}_{2}(w)=0}$ (20)
with the elliptic $r$-matrix from (16). Then the classical transfer-matrix ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)$ is a generating function of commuting Hamiltonians since
$\begin{array}[]{c}\displaystyle{\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(w)\}=0\,.}\end{array}$ (21)
Notice that $({{\mathcal{L}}}^{a})^{-1}(-z)={{\mathcal{L}}}^{a}(z)/\det({{\mathcal{L}}}^{a}(z))$ since the functions $\varphi_{k}(z)$ (12) are odd in $z$. On the other hand, $\det{{\mathcal{L}}}^{a}(z)={\bf C}_{2}^{a}-\wp(z){\bf C}_{1}^{a}=1$ due to (17). Therefore, ${\mathcal{T}}^{-1}(-z)=(-1)^{n}{\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.$ Finally,
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z){\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)K^{-}(z){\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.}\end{array}$ (22)
5 Gauge transformed $K$-matrices
Plugging the factorized representation ${\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})$ (11) into (22) we get
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z)g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\dots g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times K^{-}(z)g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\dots g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\,.}\end{array}$ (23)
Consider the gauge transformed monodromy matrix ${\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=g^{-1}(z,{\bf{q}}_{1}){\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)g(z,{\bf{q}}_{1})$:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)={\tilde{K}}^{+}(z)e^{P^{1}/2c}L^{[1,2]}(z)\dots L^{[n-1,n]}(z){\tilde{K}}^{-}(z)e^{P^{n}/2c}L^{[n,n-1]}(z)\dots L^{[2,1]}(z)\,,}\end{array}$ (24)
where the notation (3) is used and
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=g^{-1}(z,{\bf{q}}_{1}){K}^{+}(z)g(z,{\bf{q}}_{1})\,,\qquad{\tilde{K}}^{-}(z)=g^{-1}(z,{\bf{q}}_{n}){K}^{-}(z)g(z,{\bf{q}}_{n})\,.}\end{array}$ (25)
The elliptic $K$-matrices solving the reflection equation (20) are known [5]. We write them in the following form:
$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu_{0}+\sum\limits_{\alpha=1}^{3}\frac{\nu^{\pm}_{\alpha}}{c_{\alpha}(\tau)}\,\frac{1}{\varphi_{\alpha}(z)}\,\sigma_{\alpha}\,,\qquad c_{k}(\tau)=\Big(\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\quad k=1,2,3\,.}\end{array}$ (26)
Equivalently,
$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu^{\pm}_{0}+\sigma_{1}\nu^{\pm}_{1}\varphi_{1}(z+\omega_{1})-\sigma_{2}\nu^{\pm}_{2}\varphi_{2}(z+\omega_{2})-\sigma_{3}\nu^{\pm}_{3}\varphi_{3}(z+\omega_{3})\,.}\end{array}$ (27)
In order to compute the gauged transformed $K$-matrices (25) one should use identities with doubled nodular parameter $2\tau$:
$\begin{array}[]{c}\displaystyle{\theta_{2}(x+y|2\tau)\theta_{2}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{3}(x|\tau)\theta_{3}(y|\tau)-\theta_{4}(x|\tau)\theta_{4}(y|\tau)\Big)\,,}\\ \\ \displaystyle{\theta_{2}(x+y|2\tau)\theta_{3}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{2}(x|\tau)\theta_{2}(y|\tau)-\theta_{1}(x|\tau)\theta_{1}(y|\tau)\Big)\,,}\\ \\ \displaystyle{\theta_{3}(x+y|2\tau)\theta_{3}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{3}(x|\tau)\theta_{3}(y|\tau)+\theta_{4}(x|\tau)\theta_{4}(y|\tau)\Big)\,.}\end{array}$ (28)
In particular, it provides $\det g(z,{\bf{q}}_{1})=-\theta_{1}(z)\theta_{1}(2{\bf{q}}_{1})$. Direct computation yields
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1},z+\omega_{k})}\\ {\varphi_{k}(-2{\bf{q}}_{1},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\end{array}$ (29)
where $\varphi_{k}(z,x+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z+x)}{\theta_{1}(z)\theta_{5-k}(x)}$ for $k=1,2,3$, that is $\varphi_{k}(z,x+\omega_{k})|_{x=0}=\varphi_{k}(z)$. The expression for ${\tilde{K}}^{-}(z)$ is the same as (29) with $\nu^{+}_{k}$ being replaced with $\nu^{-}_{k}$ and ${\bf{q}}_{1}$ being replaced with ${\bf{q}}_{n}$.
6 Open elliptic Toda chain with boundaries
By construction,
$\begin{array}[]{c}\displaystyle{\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(w)\}=\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(w)\}=0\,.}\end{array}$ (30)
Let us compute the Hamiltonian for the open elliptic Toda chain generated by ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (24). Namely, consider the coefficient behind $1/z^{2n-2}$ in the expansion of ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ near $z=0$:
$\begin{array}[]{c}\displaystyle{H=\mathop{\hbox{Res}}\limits\limits_{z=0}z^{2n-1}{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)\,.}\end{array}$ (31)
Notice that matrices $L^{[a,b]}(z)$ (4) have simple poles at $z=0$, while the $K$-matrices ${\tilde{K}}^{\pm}(z)$ (29) are regular at $z=0$. Therefore,
$\begin{array}[]{c}\displaystyle{H={\rm tr}\Big({\tilde{K}}^{+}(0)e^{P^{1}/2c}A^{[1,2]}\dots A^{[n-1,n]}{\tilde{K}}^{-}(0)e^{P^{n}/2c}A^{[n,n-1]}\dots A^{[2,1]}\Big)\,,\quad A^{[k,m]}=\mathop{\hbox{Res}}\limits\limits_{z=0}L^{[k,m]}(z)\,.}\end{array}$ (32)
The residues $A^{[k,m]}$ are easily calculated using the property $\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1$:
$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(b_{1}^{[k,m]}\ b_{2}^{[k,m]})\,.}\end{array}$ (33)
It is also useful to mention that
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)-{\tilde{f}}^{+}({\bf{q}}_{1})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{+}({\bf{q}}_{1})=\sum\limits_{k=1}^{3}\nu^{+}_{k}\varphi_{k}(2{\bf{q}}_{1})}\end{array}$ (34)
and similarly
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{-}_{0}}\end{array}\right)-{\tilde{f}}^{-}({\bf{q}}_{n})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{-}({\bf{q}}_{n})=\sum\limits_{k=1}^{3}\nu^{-}_{k}\varphi_{k}(2{\bf{q}}_{n})\,.}\end{array}$ (35)
In particular, this yields
$\begin{array}[]{c}\displaystyle{(b_{1}^{[2,1]}\ b_{2}^{[2,1]}){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=2\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big)\Big(2\nu_{0}^{+}\cosh\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)}\end{array}$ (36)
One may use an alternatively expression for ${\tilde{f}}^{+}({\bf{q}}_{1})$ (and similarly for ${\tilde{f}}^{-}({\bf{q}}_{n})$) coming from identity
$\begin{array}[]{c}\displaystyle{\big({\tilde{f}}^{+}({\bf{q}}_{1})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\,,}\end{array}$ (37)
where
[…]
Finally, we come to the following Hamiltonian:
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\log\Big(2\nu_{0}^{+}\cosh({\bf{p}}_{1}/2c)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)+}\\ \\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\log\Big(2\nu_{0}^{-}\cosh({\bf{p}}_{n}/2c)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)+}\\ \\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (39)
Consider two particular cases.
Example 1: pure open chain. Let $\nu^{\pm}_{k}=0$, $k=1,2,3$ and $\nu^{\pm}_{0}=1$. Then we come to a ”pure” open elliptic Toda chain (corresponding to $K^{\pm}(z)=\sigma_{0}$ in XYZ chain):
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{pure\ open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)\cosh({\bf{p}}_{1}/2c)}+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)\cosh({\bf{p}}_{n}/2c)}+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (40)
Example 2: boundary terms. Let $\nu^{\pm}_{0}=0$. Then
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{boundary}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp(q_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp(q_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (41)
In conclusion, we note again that the Ruijsenaars-Toda chain [2] generalizes the elliptic Toda chain (1), and it is gauge equivalent to a generic XYZ chain (without restrictions (17)). Similarly to the results of the present paper it is then possible to derive the open Ruijsenaars-Toda chain with boundaries. The boundary $K$-matrices can be also chosen to be dynamical, see the Zhukovsky-Volterra gyrostat [7] and the 4-constant ${\rm BC}_{1}$ Ruijsenaars-van Diejen model [8]. These generalizations will be studied elsewhere.
Acknowledgments.
This work was supported by the Russian Science Foundation under grant no. 25-11-00081, https://rscf.ru/project/25-11-00081/ .

### v3

@@
-Integrable open elliptic Toda chain with boundaries
+arXiv:2602.13903v3[nlin.SI] 28 Feb 2026
 Abstract
@@
 This is an intertwining matrix entering the IRF-Vertex correspondence [4] in quantum statistical models.
+Its geometric interpretation and applications to interrelations between classical integrable systems can be found in [9, 18, 3].
 We use the standard notations for Jacobi theta functions.
@@
 Let us introduce a notation for a set of $2\times 2$ matrices:
-$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=-g^{-1}(z,{\bf{q}}_{k})g(z,{\bf{q}}_{m})e^{P^{m}/2c}\,,\qquad P^{m}={\rm diag}({\bf{p}}_{m},-{\bf{p}}_{m})}\end{array}$ (3)
-Direct calculation yields expression for $L^{[k,m]}(z)$ in terms of the elliptic Kronecker function $\phi(z,x)$:
-$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\\ {\phi(z,-{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,-{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\end{array}\right)\,,\quad\phi(z,u)=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+u)}{\theta_{1}(z)\theta_{1}(u)}}\end{array}$ (4)
-where
-$\begin{array}[]{c}\displaystyle{b^{[k,m]}_{1}=\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{{{\bf{p}}_{m}}/{2c}}\ \ b^{[k,m]}_{2}=-\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{-{{\bf{p}}_{m}}/{2c}}\,.}\end{array}$ (5)
-The Lax matrices for the elliptic Toda chain can be written in several different ways. First possibility is to define the monodromy matrix $T^{\hbox{\tiny{eToda}}}(z)$ as
-$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=L^{[n,1]}(z)L^{[1,2]}(z)L^{[2,3]}(z)\dots L^{[n-1,n]}(z)\,,}\end{array}$ (6)
-where we use the Lax matrices (4). Another possibility is to use the modified Lax representation [16, 10]:
-$\begin{array}[]{c}\displaystyle{{\bf T}^{\hbox{\tiny{eToda}}}(z)={\bf L}^{1}(z){\bf L}^{2}(z)\dots\bf L^{n}(z)\,,\qquad{\bf L}^{a}(z)=\frac{1}{b^{[a-1,a]}_{1}+b^{[a-1,a]}_{2}}\,L^{[a-1,a]}(z)\,.}\end{array}$ (7)
-Then ${\bf T}^{\hbox{\tiny{eToda}}}(z)={T}^{\hbox{\tiny{eToda}}}(z)e^{-H^{\hbox{\tiny{eToda}}}/2}$. Explicitly,
-$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{2}\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\\ {\phi(z,-{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,-{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\end{array}\right)\,.}\end{array}$ (8)
-Using the identity $\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)$ we have
-$\begin{array}[]{c}\displaystyle{\det{\bf L}^{a}(z)=\frac{1}{4}\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (9)
-Then it is easy to see that $\log\det{\bf T}^{\hbox{\tiny{eToda}}}(z)$ provides the Hamiltonian (1).
+$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=-g^{-1}(z,{\bf{q}}_{k})g(z,{\bf{q}}_{m})e^{P^{m}/2c}\,,\quad k,m=1,...,n\,,}\end{array}$ (3)
+where $P^{m}$ are diagonal $2\times 2$ matrices of particles momenta:
+$\begin{array}[]{c}\displaystyle{P^{m}={\rm diag}({\bf{p}}_{m},-{\bf{p}}_{m})\,,\quad m=1,...,n\,.}\end{array}$ (4)
+Direct calculation yields expression for $L^{[k,m]}(z)$
+$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\\ \\ {\phi(z,-{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,-{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\end{array}\right)\,,}\end{array}$ (5)
+where $\phi(z,x)$ is the elliptic Kronecker function
+$\begin{array}[]{c}\displaystyle{\phi(z,u)=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+u)}{\theta_{1}(z)\theta_{1}(u)}}\end{array}$ (6)
+and $b^{[k,m]}_{1}$, $b^{[k,m]}_{2}$ are the following functions:
+$\begin{array}[]{c}\displaystyle{b^{[k,m]}_{1}=\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{{{\bf{p}}_{m}}/{2c}}\,,}\\ \\ \displaystyle{b^{[k,m]}_{2}=-\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{-{{\bf{p}}_{m}}/{2c}}\,.}\end{array}$ (7)
+The Lax matrices for the elliptic Toda chain can be written in several different ways.
+First possibility is to use $L^{[a-1,a]}(z)$, a=1,,na=1,...,n from the definition (5).
+The monodromy matrix $T^{\hbox{\tiny{eToda}}}(z)$ takes the form [13]:
+$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=L^{[n,1]}(z)L^{[1,2]}(z)L^{[2,3]}(z)\dots L^{[n-1,n]}(z)\,.}\end{array}$ (8)
+Here $z$ is the spectral parameter.
+Being written in this from, the elliptic Toda chain is a particular case of ${\rm GL}_{N}$ elliptic Ruijsenaars chain on $n$ sites introduced in [20].
+Its classical $r$-matrix structure was described in [12].
+More precisely, ${\rm GL}_{2}$ elliptic Ruijsenaars chain depends on the additional deformation parameter $\eta$, and the elliptic Toda chain arises when $\eta=0$, while for $\eta\neq 0$ it reproduces another known integrable model – the elliptic Ruijsenaars-Toda chain [^1] In fact, the elliptic Ruijsenaars-Toda chain contains $n$ parameters $\eta_{a}$, a=1,,na=1,...,n.
+That is, the elliptic Ruijsenaars-Toda chain with $\eta_{a}=0$ for all $a$ yields the elliptic Toda chain, and the elliptic Ruijsenaars-Toda chain with pairwise equal $\eta_{a}=\eta_{b}$ corresponds to ${\rm GL}_{2}$ elliptic Ruijsenaars chain (with reduction to the center of mass frames at each site).
+Details of interrelations between models can be found in [13].
+described in [2].
+Alternatively, one can use the modified Lax representation:
+$\begin{array}[]{c}\displaystyle{{\bf T}^{\hbox{\tiny{eToda}}}(z)={\bf L}^{1}(z){\bf L}^{2}(z)\dots\bf L^{n}(z)}\end{array}$ (9)
+with
+$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{b^{[a-1,a]}_{1}+b^{[a-1,a]}_{2}}\,L^{[a-1,a]}(z)\,.}\end{array}$ (10)
+Explicitly, we have
+$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{2}\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\\ \\ {\phi(z,-{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,-{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\end{array}\right)\,.}\end{array}$ (11)
+The classical transfer-matrices in these two descriptions are related as ${\bf T}^{\hbox{\tiny{eToda}}}(z)={T}^{\hbox{\tiny{eToda}}}(z)e^{-H^{\hbox{\tiny{eToda}}}/2}$. Using the identity
+$\begin{array}[]{c}\displaystyle{\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)}\end{array}$ (12)
+for the function (6) we have
+$\begin{array}[]{c}\displaystyle{\det{\bf L}^{a}(z)=\frac{1}{4}\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)}\end{array}$ (13)
+for ${\bf L}^{a}(z)$ (11). Notice that it is independent of spectral parameter $z$. Then it is easy to see that the expression
+$\begin{array}[]{c}\displaystyle{\log\det{\bf T}^{\hbox{\tiny{eToda}}}(z)=\sum\limits_{a=1}^{n}\log\det{\bf L}^{a}(z)}\end{array}$ (14)
+provides the original Hamiltonian (1).
+Details of the above results can be found in [20, 13].
+Let us also mention that we use the factorization formulae of type (3).
+Originally, this type representation of the Lax matrices was observed by K.
+Hasegawa in [6] at quantum level for the Ruijsenaars-Schneider model.
+See also a review [18] and references therein.
 3 Gauge equivalence with XYZ chain
 Plugging the factorized representation (3) into (8) gives
-$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=(-1)^{n}g^{-1}(z,{\bf{q}}_{n})g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})g(z,{\bf{q}}_{2})e^{P^{2}/2c}...g^{-1}(z,{\bf{q}}_{n-1})g(z,{\bf{q}}_{n})e^{P^{n}/2c}}\end{array}$ (10)
-Therefore, the gauge transformed monodromy matrix ${\mathcal{T}}(z)=g(z,{\bf{q}}_{n})T^{\hbox{\tiny{eToda}}}(z)g^{-1}(z,{\bf{q}}_{n})$ takes the form
-$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=(-1)^{n}{\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)\,,\qquad{\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})}\end{array}$ (11)
-It was explained in [10] that ${\mathcal{L}}^{a}(z)$ are the Lax matrices of the XYZ chain, and ${\mathcal{T}}(z)$ is its standard monodromy matrix [12, 4]. Namely,
-$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\,,\qquad\varphi_{k}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z)}{\theta_{1}(z)\theta_{5-k}(0)}\,,\quad k=1,2,3\,,}\end{array}$ (12)
-where $\sigma_{a}$ are the Pauli matrices ($\sigma_{0}$ is $2\times 2$ identity matrix). The variables ${\bf S}^{a}_{k}$ are expressed through ${\bf{p}}_{a}$ and ${\bf{q}}_{a}$ as
-$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}=\frac{1}{2}\Big(e^{{\bf{p}}_{a}/2c}+e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad{\bf S}_{k}^{a}=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)}\end{array}$ (13)
-for all a=1,,na=1,...,n and $k=1,2,3$. The Poisson brackets between ${\bf S}_{k}^{a}$ computed through $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$ reproduce the classical Sklyanin algebra [11]
-$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i},{\bf S}^{a}_{j}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}{\bf S}^{a}_{k}\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0},{\bf S}^{a}_{i}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}{\bf S}^{a}_{k}\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,,}\end{array}$ (14)
-where the half-periods $\omega_{j}$
+$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=}\\ \\ \displaystyle{=(-1)^{n}g^{-1}(z,{\bf{q}}_{n})g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})g(z,{\bf{q}}_{2})e^{P^{2}/2c}...g^{-1}(z,{\bf{q}}_{n-1})g(z,{\bf{q}}_{n})e^{P^{n}/2c}\,.}\end{array}$ (15)
+Therefore, the gauge transformed monodromy matrix
+$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=g(z,{\bf{q}}_{n})T^{\hbox{\tiny{eToda}}}(z)g^{-1}(z,{\bf{q}}_{n})}\end{array}$ (16)
+takes the form
+$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=(-1)^{n}{\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)}\end{array}$ (17)
+with the Lax matrices
+$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=g(z,{\bf{q}}_{a})e^{P^{a}/2c}g^{-1}(z,{\bf{q}}_{a})\,,\quad a=1,...,n\,.}\end{array}$ (18)
+It was explained in [13] that ${\mathcal{L}}^{a}(z)$ are the Lax matrices of the XYZ chain, and ${\mathcal{T}}(z)$ is its standard monodromy matrix [15, 5].
+Namely, the matrices ${\mathcal{L}}^{a}(z)$ are represented in the form:
+$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\,,}\end{array}$ (19)
+where $\sigma_{a}$ are the Pauli matrices
+$\begin{array}[]{c}\displaystyle{\sigma_{0}=\left(\begin{array}[]{cc}{1}&{0}\\ {0}&{1}\end{array}\right)\,,\quad\sigma_{1}=\left(\begin{array}[]{cc}{0}&{1}\\ {1}&{0}\end{array}\right)\,,\quad\sigma_{2}=\left(\begin{array}[]{cc}{0}&{-\imath}\\ {\imath}&{0}\end{array}\right)\,,\quad\sigma_{3}=\left(\begin{array}[]{cc}{1}&{0}\\ {0}&{-1}\end{array}\right)}\end{array}$ (20)
+and
+$\begin{array}[]{c}\displaystyle{\varphi_{1}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{4}(z)}{\theta_{1}(z)\theta_{4}(0)}\,,\qquad\varphi_{2}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{3}(z)}{\theta_{1}(z)\theta_{3}(0)}\,,\qquad\varphi_{3}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{2}(z)}{\theta_{1}(z)\theta_{2}(0)}\,,}\end{array}$ (21)
+or, in the short form
+$\begin{array}[]{c}\displaystyle{\varphi_{k}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z)}{\theta_{1}(z)\theta_{5-k}(0)}\,,\quad k=1,2,3\,.}\end{array}$ (22)
+The variables ${\bf S}^{a}_{k}$ entering (19) are expressed through ${\bf{p}}_{a}$ and ${\bf{q}}_{a}$ as follows:
+$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}=\frac{1}{2}\Big(e^{{\bf{p}}_{a}/2c}+e^{-{\bf{p}}_{a}/2c}\Big)\,,}\\ \\ \displaystyle{{\bf S}_{k}^{a}=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad k=1,2,3}\end{array}$ (23)
+for all a=1,,na=1,...,n.
+This is a particular case of a more general formulae [^2] In the continuous ($n\to\infty$) and non-relativistic ($c\to\infty$) limit the XYZ chain turns into 1+1 XYZ Landau-Lifshitz model.
+It is gauge equivalent to 1+1 field analogue of 2-body Calogero-Moser model [9, 3].
+In this respect, the elliptic Toda chain in the continuous non-relativistic limit provides this 1+1 field analogue of 2-body Calogero-Moser model with zero coupling constant.:
+$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z,\eta)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}(\eta)\,,}\end{array}$ (24)
+$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}(\eta)=\frac{1}{2}\Big(\frac{\theta_{1}(2{\bf{q}}_{a}-\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}+\frac{\theta_{1}(2{\bf{q}}_{a}+\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)\,,}\\ \\ \displaystyle{{\bf S}_{k}^{a}(\eta)=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a}-\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a}+\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad k=1,2,3\,.}\end{array}$ (25)
+The Poisson brackets between ${\bf S}_{k}^{a}(\eta)$ computed through $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$ reproduce the classical Sklyanin algebra [14]:
+$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i}(\eta),{\bf S}^{a}_{j}(\eta)\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}(\eta){\bf S}^{a}_{k}(\eta)\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0}(\eta),{\bf S}^{a}_{i}(\eta)\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}(\eta){\bf S}^{a}_{k}(\eta)\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,,}\end{array}$ (26)
+where the half-periods $\omega_{j}$ (of the underlying elliptic curve with moduli $\tau$)
 $\begin{array}[]{c}\displaystyle{\omega_{1}=\frac{\tau}{2}\,,\quad\omega_{2}=\frac{1+\tau}{2}\,,\quad\omega_{3}=\frac{1}{2}}\end{array}$ (27)
-are numerated according to numeration of the Pauli matrices. The Sklyanin algebra (14) is generated by the classical quadratic exchange relation
-$\begin{array}[]{c}\displaystyle{\{{\mathcal{L}}_{1}^{a}(z),{\mathcal{L}}_{2}^{b}(w)\}=\frac{\delta^{ab}}{c}\,[{\mathcal{L}}_{1}^{a}(z){\mathcal{L}}_{2}^{a}(w),r_{12}(z-w)]\,,\qquad r_{12}(z-w)=\frac{1}{2}\sum\limits_{k=1}^{3}\varphi_{k}(z-w)\sigma_{k}\otimes\sigma_{k}\,.}\end{array}$ (16)
-The representation (13) for the variables ${\bf S}^{a}_{k}$ corresponds to special values of the Casimir functions ${\bf C}_{1}^{a}=({\bf S}^{a}_{1})^{2}+({\bf S}^{a}_{2})^{2}+({\bf S}^{a}_{3})^{2}$ and ${\bf C}_{2}^{a}=({\bf S}^{a}_{0})^{2}+\sum\limits_{k=1}^{3}({\bf S}^{a}_{k})^{2}\wp(\omega_{k})$:
+are numerated according to numeration of the Pauli matrices.
+The expressions ${\bf S}^{a}_{k}(\eta)$ (25) are the classical analogues of the representation of quantum Sklyanin algebra through difference operators (the shift operators become exponents of momenta in classical mechanics).
+The limit $\eta\to 0$ is regular for the functions ${\bf S}^{a}_{k}(\eta)$ (25), and the structure constants of the Sklyanin algebra (26) are independent of $\eta$.
+Therefore, expressions ${\bf S}^{a}_{k}$ (23) have the same Poisson brackets:
+$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i},{\bf S}^{a}_{j}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}{\bf S}^{a}_{k}\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0},{\bf S}^{a}_{i}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}{\bf S}^{a}_{k}\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,.}\end{array}$ (28)
+See details in [13] and [20].
+The Sklyanin algebra (26) is generated by the classical quadratic exchange relation for the Lax matrices (19)
+$\begin{array}[]{c}\displaystyle{\{{\mathcal{L}}_{1}^{a}(z,\eta),{\mathcal{L}}_{2}^{b}(w,\eta)\}=\frac{\delta^{ab}}{c}\,[{\mathcal{L}}_{1}^{a}(z,\eta){\mathcal{L}}_{2}^{a}(w,\eta),r_{12}(z-w)]}\end{array}$ (29)
+with the elliptic $r$-matrix
+$\begin{array}[]{c}\displaystyle{r_{12}(z-w)=\frac{1}{2}\sum\limits_{k=1}^{3}\varphi_{k}(z-w)\sigma_{k}\otimes\sigma_{k}\,.}\end{array}$ (30)
+The center of the algebra (29) (and therefore, center of (25) or (26)) is generated by
+$\begin{array}[]{c}\displaystyle{\det{{\mathcal{L}}}^{a}(z,\eta)={\bf C}_{2}^{a}(\eta)-\wp(z){\bf C}_{1}^{a}(\eta)}\end{array}$ (31)
+with the Casimir functions
+$\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}(\eta)=\big({\bf S}^{a}_{1}(\eta)\big)^{2}+\big({\bf S}^{a}_{2}(\eta)\big)^{2}+\big({\bf S}^{a}_{3}(\eta)\big)^{2}\,,}\\ \\ \displaystyle{{\bf C}_{2}^{a}(\eta)=\big({\bf S}^{a}_{0}(\eta)\big)^{2}+\sum\limits_{k=1}^{3}\big({\bf S}^{a}_{k}(\eta)\big)^{2}\wp(\omega_{k})\,.}\end{array}$ (32)
+The representation (25) yields
+$\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}(\eta)=\Big(\frac{\theta_{1}(\eta)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\qquad{\bf C}_{2}^{a}(\eta)=\wp(\eta)\Big(\frac{\theta_{1}(\eta)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,.}\end{array}$ (33)
+The representation (23) appears from (25) in the limit $\eta\to 0$.
+Then for ${\bf C}_{1}^{a}=({\bf S}^{a}_{1})^{2}+({\bf S}^{a}_{2})^{2}+({\bf S}^{a}_{3})^{2}$ and ${\bf C}_{2}^{a}=({\bf S}^{a}_{0})^{2}+\sum\limits_{k=1}^{3}({\bf S}^{a}_{k})^{2}\wp(\omega_{k})$ defined through the functions (23) we have
 $\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}=0\,,\qquad{\bf C}_{2}^{a}=1\,.}\end{array}$ (34)
+Then from (31) we also conclude that for the Lax matrices ${{\mathcal{L}}}^{a}(z)$ (19)
+$\begin{array}[]{c}\displaystyle{\det{{\mathcal{L}}}^{a}(z)=\det{{\mathcal{L}}}^{a}(z,\eta)\big|_{\eta=0}=1\,.}\end{array}$ (35)
 4 XYZ chain with boundaries
-Following [13] consider the monodromy matrix:
+Let us recall main steps for construction of integrable chains with boundaries following [16]. Consider the monodromy matrix:
 $\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z){\mathcal{T}}(z)K^{-}(z){\mathcal{T}}^{-1}(-z)\,,}\end{array}$ (36)
@@
 $\begin{array}[]{c}\displaystyle{c\{{\mathcal{T}}_{1}(z),{\mathcal{T}}_{2}(w)\}=[{\mathcal{T}}_{1}(z){\mathcal{T}}_{2}(w),r_{12}(z-w)]}\end{array}$ (37)
-due to (16), and $K^{\pm}(z)$ are $2\times 2$ matrices solving the classical quadratic reflection equation
+due to each Lax matrix ${\mathcal{L}}^{a}(z)$ satisfies the same quadratic relation (29), and $K^{\pm}(z)$ are $2\times 2$ matrices solving the classical quadratic reflection equation
 ${[K^{\pm}_{1}(z)K^{\pm}_{2}(w),r_{12}(z-w)]+\,K^{\pm}_{2}(w)r_{12}(z+w)K^{\pm}_{1}(z)-\,K^{\pm}_{1}(z)r_{12}(z+w)K^{\pm}_{2}(w)=0}$ (38)
-with the elliptic $r$-matrix from (16). Then the classical transfer-matrix ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)$ is a generating function of commuting Hamiltonians since
+with the same elliptic $r$-matrix from (29). Then the classical transfer-matrix ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)$ is a generating function of Poisson commuting Hamiltonians since
 $\begin{array}[]{c}\displaystyle{\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(w)\}=0\,.}\end{array}$ (39)
-Notice that $({{\mathcal{L}}}^{a})^{-1}(-z)={{\mathcal{L}}}^{a}(z)/\det({{\mathcal{L}}}^{a}(z))$ since the functions $\varphi_{k}(z)$ (12) are odd in $z$.
-On the other hand, $\det{{\mathcal{L}}}^{a}(z)={\bf C}_{2}^{a}-\wp(z){\bf C}_{1}^{a}=1$ due to (17).
-Therefore, ${\mathcal{T}}^{-1}(-z)=(-1)^{n}{\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.$ Finally,
+In order to compute ${\mathcal{T}}^{-1}(-z)=(-1)^{n}({\mathcal{L}}^{n}(-z))^{-1}({\mathcal{L}}^{n-1}(-z))^{-1}\dots({\mathcal{L}}^{1}(-z))^{-1}$ (17) entering (36) we need to find $({{\mathcal{L}}}^{a}(-z))^{-1}$.
+For any matrix of the form (19) we have
+$\begin{array}[]{c}\displaystyle{\Big({{\mathcal{L}}}^{a}(z)\Big)^{-1}=\frac{1}{\det({{\mathcal{L}}}^{a}(z))}\Big(\sigma_{0}{\bf S}^{a}_{0}-\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\Big)\,.}\end{array}$ (40)
+The functions $\varphi_{k}(z)$ (21) are odd in $z$:
+$\begin{array}[]{c}\displaystyle{\varphi_{k}(-z)=-\varphi_{k}(z)\,.}\end{array}$ (41)
+Therefore,
+$\begin{array}[]{c}\displaystyle{\Big({{\mathcal{L}}}^{a}(-z)\Big)^{-1}=\frac{1}{\det({{\mathcal{L}}}^{a}(-z))}\Big(\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\Big)={{\mathcal{L}}}^{a}(z)\,,}\end{array}$ (42)
+where we also used $\det({{\mathcal{L}}}^{a}(-z))=\det({{\mathcal{L}}}^{a}(z))=1$ because of (35) in our case. Thus,
+$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{-1}(-z)=(-1)^{n}{\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,,}\end{array}$ (43)
+and
 $\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z){\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)K^{-}(z){\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.}\end{array}$ (44)
@@
 Plugging the factorized representation ${\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})$ (17) into (44) we get
-$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z)g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\dots g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times K^{-}(z)g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\dots g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\,.}\end{array}$ (23)
-Consider the gauge transformed monodromy matrix ${\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=g^{-1}(z,{\bf{q}}_{1}){\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)g(z,{\bf{q}}_{1})$:
-$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)={\tilde{K}}^{+}(z)e^{P^{1}/2c}L^{[1,2]}(z)\dots L^{[n-1,n]}(z){\tilde{K}}^{-}(z)e^{P^{n}/2c}L^{[n,n-1]}(z)\dots L^{[2,1]}(z)\,,}\end{array}$ (24)
-where the notation (3) is used and
-$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=g^{-1}(z,{\bf{q}}_{1}){K}^{+}(z)g(z,{\bf{q}}_{1})\,,\qquad{\tilde{K}}^{-}(z)=g^{-1}(z,{\bf{q}}_{n}){K}^{-}(z)g(z,{\bf{q}}_{n})\,.}\end{array}$ (25)
-The elliptic $K$-matrices solving the reflection equation (20) are known [5]. We write them in the following form:
-$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu_{0}+\sum\limits_{\alpha=1}^{3}\frac{\nu^{\pm}_{\alpha}}{c_{\alpha}(\tau)}\,\frac{1}{\varphi_{\alpha}(z)}\,\sigma_{\alpha}\,,\qquad c_{k}(\tau)=\Big(\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\quad k=1,2,3\,.}\end{array}$ (26)
-Equivalently,
+$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=}\\ \\ \displaystyle{=K^{+}(z)g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\dots g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times K^{-}(z)g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\dots g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\,.}\end{array}$ (45)
+Consider the gauge transformed monodromy matrix:
+$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=g^{-1}(z,{\bf{q}}_{1}){\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)g(z,{\bf{q}}_{1})\,.}\end{array}$ (46)
+Using the notation (3) it is written in the form:
+$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=}\\ \\ \displaystyle{={\tilde{K}}^{+}(z)e^{P^{1}/2c}L^{[1,2]}(z)\dots L^{[n-1,n]}(z){\tilde{K}}^{-}(z)e^{P^{n}/2c}L^{[n,n-1]}(z)\dots L^{[2,1]}(z)\,,}\end{array}$ (47)
+where
+$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=g^{-1}(z,{\bf{q}}_{1}){K}^{+}(z)g(z,{\bf{q}}_{1})\,,}\\ \\ \displaystyle{{\tilde{K}}^{-}(z)=g^{-1}(z,{\bf{q}}_{n}){K}^{-}(z)g(z,{\bf{q}}_{n})\,.}\end{array}$ (48)
+The elliptic $K$-matrix solving the reflection equation (38) is known from [7]. We write it in the following form:
+$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu_{0}^{\pm}+\sum\limits_{\alpha=1}^{3}\frac{\nu^{\pm}_{\alpha}}{c_{\alpha}(\tau)}\,\frac{1}{\varphi_{\alpha}(z)}\,\sigma_{\alpha}}\end{array}$ (49)
+with the set of theta-constants
+$\begin{array}[]{c}\displaystyle{c_{k}(\tau)=\Big(\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\quad k=1,2,3\,.}\end{array}$ (50)
+Each $K$-matrix $K^{+}(z)$ and $K^{-}(z)$ depends on four coupling constants $\nu_{i}^{+}$ and $\nu^{-}_{i}$, $i=0,1,2,3$ respectively. Equivalently, expressions (49) are represented as follows:
 $\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu^{\pm}_{0}+\sigma_{1}\nu^{\pm}_{1}\varphi_{1}(z+\omega_{1})-\sigma_{2}\nu^{\pm}_{2}\varphi_{2}(z+\omega_{2})-\sigma_{3}\nu^{\pm}_{3}\varphi_{3}(z+\omega_{3})\,.}\end{array}$ (51)
-In order to compute the gauged transformed $K$-matrices (25) one should use identities with doubled nodular parameter $2\tau$:
+Proposition The gauge transformed $K$-matrices ${\tilde{K}}(z)$ (48) defined through $K^{\pm}(z)$ (49) with the gauge transformation (2) take the following form:
+$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ \\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1},z+\omega_{k})}\\ \\ {\varphi_{k}(-2{\bf{q}}_{1},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\\ \\ \displaystyle{{\tilde{K}}^{-}(z)=\left(\begin{array}[]{cc}{\nu^{-}_{0}}&{0}\\ \\ {0}&{\nu^{-}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{-}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{n})}&{\varphi_{k}(2{\bf{q}}_{n},z+\omega_{k})}\\ \\ {\varphi_{k}(-2{\bf{q}}_{n},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{n})}\end{array}\right)\,,}\end{array}$ (52)
+where the set of functions
+$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,x+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z+x)}{\theta_{1}(z)\theta_{5-k}(x)}\,,\quad k=1,2,3\,,}\end{array}$ (53)
+$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,x+\omega_{k})\big|_{x=0}=\varphi_{k}(z)\,.}\end{array}$ (54)
+is used.
+The proof is by direct computation using identities with doubled modular parameter $2\tau$:
 $\begin{array}[]{c}\displaystyle{\theta_{2}(x+y|2\tau)\theta_{2}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{3}(x|\tau)\theta_{3}(y|\tau)-\theta_{4}(x|\tau)\theta_{4}(y|\tau)\Big)\,,}\\ \\ \displaystyle{\theta_{2}(x+y|2\tau)\theta_{3}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{2}(x|\tau)\theta_{2}(y|\tau)-\theta_{1}(x|\tau)\theta_{1}(y|\tau)\Big)\,,}\\ \\ \displaystyle{\theta_{3}(x+y|2\tau)\theta_{3}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{3}(x|\tau)\theta_{3}(y|\tau)+\theta_{4}(x|\tau)\theta_{4}(y|\tau)\Big)\,.}\end{array}$ (55)
-In particular, it provides $\det g(z,{\bf{q}}_{1})=-\theta_{1}(z)\theta_{1}(2{\bf{q}}_{1})$. Direct computation yields
-$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1},z+\omega_{k})}\\ {\varphi_{k}(-2{\bf{q}}_{1},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\end{array}$ (29)
-where $\varphi_{k}(z,x+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z+x)}{\theta_{1}(z)\theta_{5-k}(x)}$ for $k=1,2,3$, that is $\varphi_{k}(z,x+\omega_{k})|_{x=0}=\varphi_{k}(z)$.
-The expression for ${\tilde{K}}^{-}(z)$ is the same as (29) with $\nu^{+}_{k}$ being replaced with $\nu^{-}_{k}$ and ${\bf{q}}_{1}$ being replaced with ${\bf{q}}_{n}$.
+In particular, it provides $\det g(z,{\bf{q}}_{a})=-\theta_{1}(z)\theta_{1}(2{\bf{q}}_{a})$.
+Let us remark that similar calculation was performed for relation between ${\rm BC}_{1}$ Calogero-Inozemtsev model and the non-relativistic Zhukovsky-Volterra gyrostat in [10] as well as for relation between ${\rm BC}_{1}$ Ruijsenaars- van Diejen model and the relativistic Zhukovsky-Volterra gyrostat in recent paper [11].
 6 Open elliptic Toda chain with boundaries
@@
 Let us compute the Hamiltonian for the open elliptic Toda chain generated by ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (47).
-Namely, consider the coefficient behind $1/z^{2n-2}$ in the expansion of ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ near $z=0$:
+Transfer-matrix.
+Consider the coefficient behind $1/z^{2n-2}$ in the expansion of ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ near $z=0$:
 $\begin{array}[]{c}\displaystyle{H=\mathop{\hbox{Res}}\limits\limits_{z=0}z^{2n-1}{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)\,.}\end{array}$ (57)
-Notice that matrices $L^{[a,b]}(z)$ (4) have simple poles at $z=0$, while the $K$-matrices ${\tilde{K}}^{\pm}(z)$ (29) are regular at $z=0$. Therefore,
-$\begin{array}[]{c}\displaystyle{H={\rm tr}\Big({\tilde{K}}^{+}(0)e^{P^{1}/2c}A^{[1,2]}\dots A^{[n-1,n]}{\tilde{K}}^{-}(0)e^{P^{n}/2c}A^{[n,n-1]}\dots A^{[2,1]}\Big)\,,\quad A^{[k,m]}=\mathop{\hbox{Res}}\limits\limits_{z=0}L^{[k,m]}(z)\,.}\end{array}$ (32)
-The residues $A^{[k,m]}$ are easily calculated using the property $\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1$:
-$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(b_{1}^{[k,m]}\ b_{2}^{[k,m]})\,.}\end{array}$ (33)
-It is also useful to mention that
-$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)-{\tilde{f}}^{+}({\bf{q}}_{1})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{+}({\bf{q}}_{1})=\sum\limits_{k=1}^{3}\nu^{+}_{k}\varphi_{k}(2{\bf{q}}_{1})}\end{array}$ (34)
+Notice that matrices $L^{[a,b]}(z)$ (5) have only simple poles at $z=0$ due to
+$\begin{array}[]{c}\displaystyle{\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1}\end{array}$ (58)
+for the function $\phi(z,u)$ (6).
+The product ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (47) contains 2n22n-2 matrices $L^{[a,b]}(z)$, and the boundary $K$-matrices ${\tilde{K}}^{\pm}(z)$ (52) are regular at $z=0$.
+Therefore,
+$\begin{array}[]{c}\displaystyle{H={\rm tr}\Big({\tilde{K}}^{+}(0)e^{P^{1}/2c}A^{[1,2]}\dots A^{[n-1,n]}{\tilde{K}}^{-}(0)e^{P^{n}/2c}A^{[n,n-1]}\dots A^{[2,1]}\Big)}\end{array}$ (59)
+with
+$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\mathop{\hbox{Res}}\limits\limits_{z=0}L^{[k,m]}(z)\,.}\end{array}$ (60)
+The residues $A^{[k,m]}$ are easily calculated using the property (58):
+$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\left(\begin{array}[]{cc}{b_{1}^{[k,m]}}&{b_{2}^{[k,m]}}\\ {b_{1}^{[k,m]}}&{b_{2}^{[k,m]}}\end{array}\right)=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(b_{1}^{[k,m]}\,,b_{2}^{[k,m]})\,,}\end{array}$ (61)
+where $(b_{1}^{[k,m]}\,,b_{2}^{[k,m]})$ are 2-dimensional row-vectors.
+Expressions for ${\tilde{K}}^{\pm}(0)$.
+Taking into account the oddness (41) of $\varphi_{k}(z)$ and the definitions (53)-(54), from (52) we have
+$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ \\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1})}\\ \\ {-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\\ \\ \displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{-}_{0}}&{0}\\ \\ {0}&{\nu^{-}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{-}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{n})}&{\varphi_{k}(2{\bf{q}}_{n})}\\ \\ {-\varphi_{k}(2{\bf{q}}_{n})}&{\varphi_{k}(2{\bf{q}}_{n})}\end{array}\right)\,,}\end{array}$ (62)
+Put it differently,
+$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)-{\tilde{f}}^{+}({\bf{q}}_{1})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\,,-1)}\end{array}$ (63)
+with
+$\begin{array}[]{c}\displaystyle{{\tilde{f}}^{+}({\bf{q}}_{1})=\sum\limits_{k=1}^{3}\nu^{+}_{k}\varphi_{k}(2{\bf{q}}_{1})}\end{array}$ (64)
 and similarly
-$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{-}_{0}}\end{array}\right)-{\tilde{f}}^{-}({\bf{q}}_{n})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{-}({\bf{q}}_{n})=\sum\limits_{k=1}^{3}\nu^{-}_{k}\varphi_{k}(2{\bf{q}}_{n})\,.}\end{array}$ (35)
-In particular, this yields
-$\begin{array}[]{c}\displaystyle{(b_{1}^{[2,1]}\ b_{2}^{[2,1]}){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=2\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big)\Big(2\nu_{0}^{+}\cosh\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)}\end{array}$ (36)
-One may use an alternatively expression for ${\tilde{f}}^{+}({\bf{q}}_{1})$ (and similarly for ${\tilde{f}}^{-}({\bf{q}}_{n})$) coming from identity
-$\begin{array}[]{c}\displaystyle{\big({\tilde{f}}^{+}({\bf{q}}_{1})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\,,}\end{array}$ (37)
+$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{-}_{0}}&{0}\\ {0}&{\nu^{-}_{0}}\end{array}\right)-{\tilde{f}}^{-}({\bf{q}}_{n})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\,,-1)\,,}\end{array}$ (65)
+$\begin{array}[]{c}\displaystyle{{\tilde{f}}^{-}({\bf{q}}_{n})=\sum\limits_{k=1}^{3}\nu^{-}_{k}\varphi_{k}(2{\bf{q}}_{n})\,.}\end{array}$ (66)
+One may use an alternative representation for the functions ${\tilde{f}}^{+}({\bf{q}}_{1})$ and ${\tilde{f}}^{-}({\bf{q}}_{n})$:
+$\begin{array}[]{c}\displaystyle{\big({\tilde{f}}^{+}({\bf{q}}_{1})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\,,}\\ \\ \displaystyle{\big({\tilde{f}}^{-}({\bf{q}}_{n})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp({\bf{q}}_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\,,}\end{array}$ (67)
 where
-[…]
-Finally, we come to the following Hamiltonian:
-$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\log\Big(2\nu_{0}^{+}\cosh({\bf{p}}_{1}/2c)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)+}\\ \\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\log\Big(2\nu_{0}^{-}\cosh({\bf{p}}_{n}/2c)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)+}\\ \\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (39)
+$\begin{array}[]{c}\left(\begin{array}[]{c}\bar{\nu}^{\pm}_{0}\\ \bar{\nu}^{\pm}_{1}\\ \bar{\nu}^{\pm}_{2}\\ \bar{\nu}^{\pm}_{3}\end{array}\right)=\displaystyle{\frac{1}{2}}\left(\begin{array}[]{ccc}1&1&1\\ 1&-1&-1\\ -1&1&-1\\ -1&-1&1\end{array}\right)\left(\begin{array}[]{c}\nu^{\pm}_{1}\\ \nu^{\pm}_{2}\\ \nu^{\pm}_{3}\end{array}\right)\,.\end{array}$ (68)
+This type identities can be found in [11].
+The Hamiltonian.
+Let us compute the expression for $H$ (59). Plugging (61) into (59) we get:
+$\begin{array}[]{c}\displaystyle{H=\prod\limits_{a=2}^{n-1}\Big(b_{1}^{[a-1,a]}+b_{2}^{[a-1,a]}\Big)\Big(b_{1}^{[a+1,a]}+b_{2}^{[a+1,a]}\Big)\times}\\ \\ \displaystyle{\times\left[\big(b_{1}^{[2,1]}\,,b_{2}^{[2,1]}\big){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)\right]\times\left[\big(b_{1}^{[n-1,n]}\,,b_{2}^{[n-1,n]}\big){\tilde{K}}^{-}(0)e^{P^{n}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)\right]\,.}\end{array}$ (69)
+It follows from (7) that
+$\begin{array}[]{c}\displaystyle{b_{1}^{[a-1,a]}+b_{2}^{[a-1,a]}=2\frac{\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{a})}\sinh\big(\frac{{\bf{p}}_{a}}{2c}\big)\,,}\\ \\ \displaystyle{b_{1}^{[a+1,a]}+b_{2}^{[a+1,a]}=2\frac{\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a+1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a+1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{a})}\sinh\big(\frac{{\bf{p}}_{a}}{2c}\big)\,.}\end{array}$ (70)
+Then for the upper line of (69) one obtains:
+$\begin{array}[]{c}\displaystyle{\prod\limits_{a=2}^{n-1}\Big(b_{1}^{[a-1,a]}+b_{2}^{[a-1,a]}\Big)\Big(b_{1}^{[a+1,a]}+b_{2}^{[a+1,a]}\Big)=}\\ \\ \displaystyle{=\theta_{1}({\bf{q}}_{2}-{\bf{q}}_{1})\theta_{1}({\bf{q}}_{2}+{\bf{q}}_{1})\theta_{1}({\bf{q}}_{n-1}-{\bf{q}}_{n})\theta_{1}({\bf{q}}_{n-1}+{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times(-1)^{n-3}\Big(\frac{2}{\theta_{1}^{\prime}(0)}\Big)^{2n-4}\,\left(\frac{\displaystyle{\prod\limits_{a=3}^{n-1}\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}}{\displaystyle{\prod\limits_{a=2}^{n-1}\theta_{1}(2{\bf{q}}_{a})}}\right)^{2}\,\prod\limits_{a=2}^{n-1}\sinh^{2}\big(\frac{{\bf{p}}_{a}}{2c}\big)\,.}\end{array}$ (71)
+Consider the lower line of (69). Using (7) and (63) we have
+$\begin{array}[]{c}\displaystyle{\big(b_{1}^{[2,1]}\,,b_{2}^{[2,1]}\big){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=}\\ \\ \displaystyle{=\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\big(e^{{\bf{p}}_{1}/2c}\,,-e^{-{\bf{p}}_{1}/2c}\big){\tilde{K}}^{+}(0)\left(\begin{array}[]{c}e^{{\bf{p}}_{1}/2c}\\ e^{-{\bf{p}}_{1}/2c}\end{array}\right)=}\\ \\ \displaystyle{=\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\Big(2\nu_{0}^{+}\sinh\big(\frac{{\bf{p}}_{1}}{c}\big)-4\sinh^{2}\big(\frac{{\bf{p}}_{1}}{2c}\big){\tilde{f}}^{+}({\bf{q}}_{1})\Big)=}\\ \\ \displaystyle{=4\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\Big(\nu_{0}^{+}\cosh\big(\frac{{\bf{p}}_{1}}{2c}\big)-\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big){\tilde{f}}^{+}({\bf{q}}_{1})\Big)=}\\ \\ \displaystyle{=4\sinh^{2}\big(\frac{{\bf{p}}_{1}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)\,.}\end{array}$ (72)
+Similarly,
+$\begin{array}[]{c}\displaystyle{\big(b_{1}^{[n-1,n]}\,,b_{2}^{[n-1,n]}\big){\tilde{K}}^{-}(0)e^{P^{n}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=}\\ \\ \displaystyle{=4\sinh\big(\frac{{\bf{p}}_{n}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{n}-{\bf{q}}_{n-1})\theta_{1}({\bf{q}}_{n}+{\bf{q}}_{n-1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{n})}\Big(\nu_{0}^{-}\cosh\big(\frac{{\bf{p}}_{n}}{2c}\big)-\sinh\big(\frac{{\bf{p}}_{n}}{2c}\big){\tilde{f}}^{-}({\bf{q}}_{n})\Big)=}\\ \\ \displaystyle{=4\sinh^{2}\big(\frac{{\bf{p}}_{n}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{n}-{\bf{q}}_{n-1})\theta_{1}({\bf{q}}_{n}+{\bf{q}}_{n-1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{n})}\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.}\end{array}$ (73)
+Plugging (71), (72) and (73) into (69) we come to the following expression for $H$:
+$\begin{array}[]{c}\displaystyle{H=(-1)^{n-1}\frac{2^{n}}{\theta_{1}^{\prime}(0)^{2n-2}}\,\frac{1}{\theta_{1}(2{\bf{q}}_{1})\theta_{1}(2{\bf{q}}_{n})}\,\left(\frac{\displaystyle{\prod\limits_{a=2}^{n}\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}}{\displaystyle{\prod\limits_{a=2}^{n-1}\theta_{1}(2{\bf{q}}_{a})}}\right)^{2}\,\prod\limits_{a=1}^{n}\sinh^{2}\big(\frac{{\bf{p}}_{a}}{2c}\big)\times}\\ \\ \displaystyle{\times\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.}\end{array}$ (74)
+Next, let us once more use the identity (12):
+$\begin{array}[]{c}\displaystyle{\wp({\bf{q}}_{a}-{\bf{q}}_{a-1})-\wp({\bf{q}}_{a}+{\bf{q}}_{a-1})=-\phi({\bf{q}}_{a}-{\bf{q}}_{a-1},{\bf{q}}_{a}+{\bf{q}}_{a-1})\phi({\bf{q}}_{a-1}-{\bf{q}}_{a},{\bf{q}}_{a}+{\bf{q}}_{a-1})=}\\ \\ \displaystyle{=-\frac{\theta_{1}^{\prime}(0)^{2}\theta_{1}(2{\bf{q}}_{a})\theta_{1}(2{\bf{q}}_{a-1})}{\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})^{2}\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})^{2}}\,.}\end{array}$ (75)
+Then
+$\begin{array}[]{c}\displaystyle{\prod\limits_{a=2}^{n}\frac{1}{\wp({\bf{q}}_{a}-{\bf{q}}_{a-1})-\wp({\bf{q}}_{a}+{\bf{q}}_{a-1})}=}\\ \\ \displaystyle{=\frac{(-1)^{n-1}}{\theta_{1}^{\prime}(0)^{2n-2}}\,\frac{1}{\theta_{1}(2{\bf{q}}_{1})\theta_{1}(2{\bf{q}}_{n})}\,\left(\frac{\displaystyle{\prod\limits_{a=2}^{n}\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}}{\displaystyle{\prod\limits_{a=2}^{n-1}\theta_{1}(2{\bf{q}}_{a})}}\right)^{2}\,.}\end{array}$ (76)
+By comparing (74) and (76) one gets
+$\begin{array}[]{c}\displaystyle{H=2^{n}\,\prod\limits_{a=2}^{n}\frac{1}{\wp({\bf{q}}_{a}-{\bf{q}}_{a-1})-\wp({\bf{q}}_{a}+{\bf{q}}_{a-1})}\,\prod\limits_{a=1}^{n}\sinh^{2}\big(\frac{{\bf{p}}_{a}}{2c}\big)\times}\\ \\ \displaystyle{\times\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.}\end{array}$ (77)
+Finally, we come to the following Hamiltonian for the open elliptic Toda chain with boundaries:
+$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{openToda}}}=-\log(H)+\log 2^{n}=}\\ \\ \displaystyle{=\sum\limits_{a=1}^{n}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)-}\\ \\ \displaystyle{-\log\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)-\log\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.\par}\end{array}$ (78)
+To summarize, we proved the following statement.
+Theorem The Hamiltonian for the open elliptic Toda chain with boundaries evaluated from the classical monodromy matrix ${\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (47) has the form (78), where the functions ${\tilde{f}}^{+}({\bf{q}}_{1})$, ${\tilde{f}}^{-}({\bf{q}}_{n})$ are defined in (64), (66) or (67).
+By construction, this model is gauge equivalent to the open XYZ chain with boundaries given by $K$-matrices (49), and with the Casimir functions fixed as in (34) at each site.
+Compared to the original Hamiltonian (1) for the closed elliptic Toda chain, the expression (78) does not contain interaction of $n$-th particle with the first one.
+But it contains boundary terms for these two particles on boundaries.
 Consider two particular cases.
-Example 1: pure open chain. Let $\nu^{\pm}_{k}=0$, $k=1,2,3$ and $\nu^{\pm}_{0}=1$. Then we come to a ”pure” open elliptic Toda chain (corresponding to $K^{\pm}(z)=\sigma_{0}$ in XYZ chain):
-$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{pure\ open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)\cosh({\bf{p}}_{1}/2c)}+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)\cosh({\bf{p}}_{n}/2c)}+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (40)
-Example 2: boundary terms. Let $\nu^{\pm}_{0}=0$. Then
-$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{boundary}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp(q_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp(q_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (41)
-In conclusion, we note again that the Ruijsenaars-Toda chain [2] generalizes the elliptic Toda chain (1), and it is gauge equivalent to a generic XYZ chain (without restrictions (17)).
+Example 1: pure open chain.
+Let $\nu^{\pm}_{k}=0$, $k=1,2,3$ and $\nu^{\pm}_{0}=1$.
+From viewpoint of the open XYZ chain this case corresponds to trivial $K$-matrices $K^{\pm}(z)=\sigma_{0}$ (49), and the gauge transformed $K$-matrices (52) are trivial as well: ${\tilde{K}}^{\pm}(z)=\sigma_{0}$.
+Then
+$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{pure\ open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)\cosh({\bf{p}}_{1}/2c)}+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)\cosh({\bf{p}}_{n}/2c)}+}\\ \\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (79)
+This Hamiltonian differs from (1) by the absence of interaction between the first and the last particles, and by type of dependence on momenta for these two particles on the boundaries.
+Example 2: only boundary terms. Now let it be vice versa: $\nu^{\pm}_{0}=0$. Using (67) we obtain
+$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{boundary}}}=\sum\limits_{a=1}^{n}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)-}\end{array}$ (80)
+${-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\Big)-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp({\bf{q}}_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\Big)\,.}$
+In this case the kinetic part (dependence on momenta) is the same as in (1) but there are additional external fields for the first and the last particles on the boundaries.
+In conclusion, we notice again that the Ruijsenaars-Toda chain [2] generalizes the elliptic Toda chain (1), and it is gauge equivalent to a generic XYZ chain (without restrictions (34)).
 Similarly to the results of the present paper it is then possible to derive the open Ruijsenaars-Toda chain with boundaries.
@@
 Acknowledgments.
-This work was supported by the Russian Science Foundation under grant no. 25-11-00081, https://rscf.ru/project/25-11-00081/ .
+We are grateful to D. Murinov and A. Zabrodin for discussions.
+This work was supported by the Russian Science Foundation under grant no. 25-11-00081,
+https://rscf.ru/project/25-11-00081/ .